Calculate the number of grams of oxygen that would react with 0.125 g Fes
1
Expert's answer
2021-10-07T03:48:26-0400

The balanced chemical equation will be

4FeS + 7O2 --> 2Fe2O3 + 4SO2


Amount of FeS given = 0.125g

Molecular Weight of FeS = 55+32 = 87

Moles of FeS = 0.125/87


Now as per the balanced equation, 4 Moles of FeS react with 7 Moles of O2

Therefore, 0.125/87 moles of FeS will react with = (7/4)*(0.125/87) = 0.0025 moles of O2


0.0025 = amount in gram / 32 => amount in gram = 0.0025*32 = 0.08g


So Amount of oxygen required is 0.08g